Solve the system of equations using the addition (elimination) method.
3x – 11y = 9
-9x + 33y = 18

Answer by Nate(3500)

We need to eliminate terms by multiplying, then adding...
(3x – 11y = 9)3
9x - 33y = 27
-9x + 33y = 18
0 = 45
There are no solutions because 0 does not equal 45!